• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:06:52 on May27,2007

    Entry number 203335

    keyword=Owl Shift Summary

    Owl Shift Summary:

    M.Olson, R. Michaels

    Shift begins with run 3782 in progress:

    0007: MCC request drop in current to 45uA to decrease trip rate.

    0132: End 3782 (He3, 45uA, 1.25M)
    End run failure for 3782
    0136: Restarting coda. ROC1 appears dead. Rebooting all ROCs
    Restart coda. Successful
    0138: Begin run 3783

    0343: End run 3783 (He3, 1.25M, 45uA), Begin run 3784

    0541: End run 3784 (He3, 1.25M, 45uA), Begin run 3785

    0747: End run 3785 (He3, 1.25M, 45uA), Begin run 3786

    0757: DAQ down, lost ROC4, end run 3786
    kcoda. rebooting all ROC's
    ROC4 did not reboot, reset from Hall A tools, crate resets.
    ROC3 and 4 rebooted successfully.
    Restarted CODA

    0804: Start run 3787